# Defect recombination origin of low energy excess in semiconductor detectors

### v1

@@ II.2 Quantum thermal bath simulations
To assess how significant quantum mechanical zero point vibrations [55] are on the annealing time scale and mechanisms, we employed the quantum thermal bath (QTB) approach [37, 36] as implemented in LAMMPS. This QTB thermostat is a Langevin-style thermostat with a specially tuned noise term that agrees with the energy spectrum of quantized vibrations in crystals. This thermostat does not cause significant extra computational cost compared with standard molecular dynamics, which is a great advantage compared with other methods, for instance, path integral molecular dynamics[56].
The equation of motion is a Langevin style formula, where $f(x)$ is interatomic force, $\gamma$ is the damping constant and $\Theta$ is the noise term:
$\ddot{x}=f(x)+\sqrt{2m\gamma}\Theta(t)-m\gamma\dot{x}$ (1)
The power spectral density of the noise term $\Theta$ is:
@@
In the implementation we use [37], the quantum thermal bath has five input parameters, namely, target quantum temperature, damping parameter, a random seed, upper cutoff frequency and the number of frequency bins. To make the results comparable with the classical molecular dynamics simulations, we set the damping parameter to be 0.3 ps (300 fs), the same one used in the purely classical simulations. The upper cutoff frequency was set to be three times the Debye frequency of silicon. The number of frequency bins was 100 in all QTB simulations. When using the thermostat, a Berendsen barostat with the settings in section II.1 was also applied to the system.
To check whether these settings enable the compensation of low temperature quantum effect, the heat capacity of pure silicon crystal was calculated by differentiating the mean total energy with respect to temperature, and the result was compared with experiment [57]. The heat capacity curve shown in Fig. 1 demonstrates that QTB thermostat can reproduce the temperature dependence of heat capacity at low temperatures.
With the help of quantum thermal bath, the annealing behavior of the system at very low temperatures can be investigated directly. The five-stage simulation and analysis procedure in section II.1 was followed, and quantum thermal bath was used in stage(ii) for the annealing simulations.
Figure 1: Heat capacity of silicon crystal
III Results and discussion
@@ III.1 Stage (i) Damage production in cascade simulations
An example of a cascade event and the damage produced is shown in a cross-sectional plot in Fig.2. The cross-section shows a very typical damage production event, analyzed in detail in many previous works [58, 28, 52]. As noted above, 5 keV is above the subcascade threshold breakdown, so the damage is split into separate damage production regions. 1000 different 5 keV cascades were simulated, and the final atom coordinates from these were used for annealing runs, quenching and energy release analysis.
Figure 2: Cross-sectional view of the damage produced in a 5 keV cascade in Si, as modelled with the Stillinger-Weber potential. The top part shows the evolution of the cascade, and the larger frame at the bottom the final damage state at 10 ps. This cascade was ran at 0 K ambient temperature to ensure there is no thermal annealing. The precise size of the plotted simulation cell region is 81.2 Å in the $x$ (horizontal), 71.3 Å in the $y$ (vertical) and 10.86 Å in the $z$ (out of plane) direction. The colors indicate the potential energy $E_{\rm pot}$ of the atoms relative to the ground state energy $E_{\rm pot,0}$
@@ III.2 Stage (ii) Annealing simulations
For the classical molecular dynamics simulations, annealing simulations were performed at several temperatures in the range 20 - 800 K for some of the cells, and at 200 K for all the cells. For the Tersoff potential, annealing was done at 300 K for all the cells (since the Tersoff potential has a higher melting point than the Stillinger-Weber and experimental one, 300 K corresponds to roughly the same homologous temperature as 200 K for the Stillinger-Weber potential). For simulations with QTB thermostat, the annealing simulations of 10 ns were conducted at 0.05 K, 10 K, 25 K, 50 K, 100 K, 150 K, 200 K and 300 K. At each temperature, at least 200 annealing simulations with different initial damage were carried out.
Results for the number of defects in the cell as a function of time are shown in Fig. 3 a) for a test 600 K annealing run via classical molecular dynamics. The potential energy has huge fluctuations because a thermodynamic system with a small number of atoms $N$ has natural fluctuations of the kinetic energy around the average of the order of $\sqrt{N}/N$ [59]. Since the total energy remains the same on short time scales, the fluctuations in kinetic energy are also reflected in the potential energy. The fluctations make it very difficult to obtain a quantitative analysis of the potential energy release in a single recombination event. This was the reason why we took into use the quenching scheme to get rid of thermal fluctuations.
@@
Figure 3: a) Annealing of damage produced in a single 5 keV cascade in Si at 600 K. The data shows that both the average potential energy and number of Wigner-Seitz defect pairs decreases with time. b) Same as a) but analyzed after each time step was quenched to 0 K. The inset shows the atomic positions in one disordered pocket before and after the annealing event occuring at $t=1.236$ ns. c) Potential energy release per 2 ps intervals from the same cascade. The inset shows that annealing at well-separated times keep occurring also at longer times and even at the cryogenic temperature of 100 K.
@@ III.3 Stage (iii) Quenching runs
Results of the potential energy and number of WS defect pairs after the quenching scheme (see section II) are shown in Fig. 3 b), for the exact same annealing run as shown in part 3 a). This graph shows much clearer that after the first about 10 ps, the annealing proceeds stepwise, with clear energy release events occuring every now and then, separated by (on the atomic scale) long times when essentially nothing else than thermal vibrations of atoms around their stable or metastable local equilibrium positions happen in the system. The energy releases also correlate very well with decreases of the number of Wigner-Seitz defects, confirming that the potential energy release is indeed associated with partial annealing of damage in the amorphous pockets.
From the potential energy release, one can also obtain the magnitude of the energy released in each annealing event by comparing the total potential energy of the system (after quenching) before and after each 2 ps annealing interval. This is shown in Fig. 3 c) for the same annealing run as in parts a) and b). The graph shows that the energy release varies from an annealing event to another.
@@
We ran similar annealing runs for a few cascades in the temperature range 20 – 800 K. Results for an exemplary case in Fig. 4 show that, as expected, at the highest temperatures efficient annealing of the damage is observed. This is in agreement with the experimental observations that Si cannot be amorphized above about 600 K [60].
However, we also observe that even at the much lower temperatures, down to 100 K, damage recombination is observed throughout the 3 ns simulations. Also in the test 30 ns simulations, recombination events kept occurring throughout the entire simulation time: an example of recombination occurring around 15 ns after the cascade event at 100 K is seen in the inset of Fig. 3 c). We note that Beland and Mousseau have earlier, at 300 K with a hybrid molecular dynamics/barrier search approach, observed annealing events on macrosopic, up to 1 s, time scales [35]. Occasional recombination events are observed in the 50 and 20 K simulations even in the classical simulations with no quantum thermal vibrations. An explanation to this surprising observation is given in section IV.
@@ III.4 Stage (iv) Statistics of energy release
To enable a comparison with the experimental observations of the energy dependence of the low-energy excess, we made a statistics of the annealing data similar to that shown in Fig. 4 but summing up all events over at least 50 independent cases (separate cascades) to obtain a statistically better picture of the annealing.
Since the main aim of the current work is to compare the results with the experimentally observed exponential low-energy excess signal, we plot the results with a logarithmic $y$ axis vs. linear $x$ axis for the energy.
For classical molecular dynamics simulations the results for the temperatures 100 - 500 K are shown in Fig. 5 (a). Also included in this plot are results obtained with the Tersoff interatomic potential at 300 K. The results show that, remarkably, the energy release statistics is independent of temperature in the range 100 - 500 K within the statistical uncertainty. This gives high confidence that also lower temperatures would show a similar dependence with energy. Moreover, the Tersoff interatomic potential gives results practically identical to those of the Stillinger-Weber potential. Since the Tersoff and Stillinger-Weber potential are developed completely independent of each other, have different functional forms, and are fitted to different properties [50, 47], this gives high confidence that the obtained results are not an interatomic potential artifact.
For simulations with QTB themostat, the results at various temperatures are shown in Fig. 5 (b). It is very clear that the energy release statistics is independent of temperature.
In both Fig. 5 (a) and Fig. 5 (b), the data are clearly close to linear on the log-lin plot, showing that the energy release statistics follows an exponential dependence with energy, in agreement with the experimental observations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. As expected from the linear dependence of the plots, the data can be fit very well with the functional form
$f(E)=Ae^{-\alpha E}$ (3)
@@
a) Refer to caption b) Refer to caption
Figure 5: Statistics of the magnitude of the energy release events during annealing runs at different temperatures. To enable easy comparison of the slopes, the curves at different temperatures have been scaled to 1 at 1 eV. a) Classical molecular dynamics, b) Molecular dynamics with quantum thermal bath
@@
The simulated and experimental data from a silicon detector are compared to each other in Fig. 6. The $y$ axis scale is in arbitrary units, since in experiments the absolute values of the signal depends on factors such as detector size, detection efficiency, concentration of radioactive impurities creating the initial damage, and damage recoil spectrum, all of which are not fully well known. However, the functional dependence of the signal on energy can be used for a direct comparison between simulation and experiments. The figure shows a very good agreement: both the simulation and low-energy tail of the experimental data clearly follows an exponential trend with a very similar slope. The slope obtained from a fit to the simulation data at 0.05 K (QTB thermostat) was $\alpha_{\rm sim}=0.58\pm 0.05$ 1/eV, which is in perfect agreement with a fit of the same exponential law to the low-energy tail of the experimental data citeHei22 $\alpha_{\rm exp}=0.57\pm 0.01$ 1/eV (to enable direct comparison, this value was obtained by fitting the experimental data with the same software as used to fit the simulated ones). We emphasize that no fitting of the simulations to the experiments has been used for obtaining this outstanding agreement.
Table 1: Slope coefficients $\alpha$ of the exponential dependence $\exp(-\alpha E)$ of the energy release $E$ from simulations and experiments on Si. The results obtained with classical molecular dynamics results are marked ”Class.” and the ones with the quantum thermal bath ”Qtb.”.
@@
Simulation 500 Class. 0.49 $\pm$ 0.03
(current work) 400 Class. 0.49 $\pm$ 0.03
300 Class. 0.53 $\pm$ 0.03
300 Qtb. 0.52 $\pm$ 0.01
200 Class. 0.52 $\pm$ 0.02
200 Qtb. 0.51 $\pm$ 0.02
100 Class. 0.60 $\pm$ 0.05
100 Qtb. 0.58 $\pm$ 0.07
50 Qtb. 0.52 $\pm$ 0.06
25 Qtb. 0.49 $\pm$ 0.04
10 Qtb. 0.50 $\pm$ 0.06
0.05 Qtb. 0.58 $\pm$ 0.05
@@ III.5 Stage (v) Analysis of annealing time scales
We finally analyze the time dependence of the annealing events such as those shown in Fig. 3 and 4. To be able to roughly estimate the annealing time scales as a function of temperature and extrapolate to even lower temperatures, we analyzed the time dependence of the annealing curves.
@@
Since the number of disordered pockets in a single event is small, the annealing curves of single events have huge statistical fluctuations. This is illustrated in Fig. 7. To obtain a statistically reliable analysis, we carried out an analysis of at least 200 different cascade events in the temperature range 0.05 – 300 K. To be able to compare cases with different numbers of initial defects and stored energy in them, the analysis was done for the average relative fraction of potential energy remaining at time $t$ compared to the initial potential energy at $t=0$. The fraction $f$ was calculated for each cascade event separately before averaging. The results of the averaged fraction analysis are shown in Fig. 8. As expected, time evolution of the fraction of remaining Wigner-Seitz defects showed a similar behaviour.
The averaged annealing curves show that the annealing behaviour can be divided in two stages: a rapid annealing occurring roughly during the first 0.5 ns after the damage event, and a subsequent much slower annealing after this. The first stage clearly is due to recombination of the most metastable defects, whereas the later recombination is an average over occasional well-separated recombination events (corresponding to the late energy release events shown in Fig. 3).
By comparing the individual cases shown in Fig. 7, an additional important point is illustrated: the damage in some events (those with large damage pockets) can anneal much slower than the average release. On long time scales, annealing from such cases would dominate the residual recovery signal.
Figure 8: Average potential energy curves after quenching (with quantum thermal bath). The values are rescaled so that the initial value is one. In this way, the curves show the relative amount of excessive potential energy.
Figure 9: Fit of annealing time parameter.
Table 2: Time parameter $\tau$ in the fitting $A\exp(-t/\tau)$ of the potential energy curves. The error bar is the uncertainty estimate of the parameter obtained from the least squares fitting procedure. The results obtained with classical molecular dynamics are marked “Class.” and the ones with the quantum thermal bath “Qtb.”
@@
150 Class. 348.8 $\pm$ 1.4
150 Qtb. 172.53 $\pm$ 0.46
125 Class. 451.5 $\pm$ 1.8
100 Class. 1037.8 $\pm$ 3.1
100 Qtb. 388.0 $\pm$ 1.3
50 Qtb. 769.8 $\pm$ 2.5
25 Qtb. 1070.9 $\pm$ 2.4
10 Qtb. 1234.0 $\pm$ 3.3
0.05 Qtb. 1326.1 $\pm$ 3.7
The curves in Fig. 8 were fitted against the exponential function $A\exp(-t/\tau)$. In the fitting practice, the initial 1 ns in the potential energy curves was truncated, and the rest of the curves were fitted. The fitting results and the comparison with the results from purely classical annealing are shown in Fig. 9 and table 2.
From these results, it is very clear that the low temperature quantum effect can facilitate annealing significantly. The simulation at 0.05 K with QTB thermostat shows that the annealing time parameter $\tau$ has a magnitude of microseconds at the operation temperature of the detectors. The fact that at 10 K and 0.05 K time constants are almost the same indicates that at temperatures $\lesssim 10$ K, the triggering of annealing effects is dominated by the quantum effect.
The exponential decay of annealing effects could be detectable if the time resolution of the experiments is shorter than the time constants $\tau$. Conversely, in case the time resolution is much longer than this, the energy release events would seem to appear at random times with average constant rate (that would be determined by the concentration of radioactive impurities that produce the initial damage).

### v2

@@ II.2 Quantum thermal bath simulations
To assess how significant quantum mechanical zero point vibrations [64] were on the annealing time scale and mechanisms, we employed the quantum thermal bath (QTB) approach [39, 38] as implemented in LAMMPS. This QTB thermostat is a Langevin-style thermostat with a specially tuned random force term that agrees with the energy spectrum of quantized vibrations in crystals. This thermostat does not incur very heavy extra computational cost compared with standard molecular dynamics, only roughly doubling the computation time of molecular dynamics simulations, which is a great advantage compared with other methods, for instance, path integral molecular dynamics [65].
When using QTB, the equation of motion is a Langevin style formula, where $f(x)$ is interatomic force, $\gamma$ is the damping constant, $t$ is time and $\Theta$ is the random force term:
$m\ddot{x}=f(x)+\sqrt{2m\gamma}\Theta(t)-m\gamma\dot{x}$ (1)
The power spectral density of the random force term $\Theta$ is:
@@
where $\omega$ is the vibration frequency, $\hbar$ is reduced Planck constant, $k_{B}$ is Boltzmann constant and T is the temperature of the system.
In the implementation we use [39], the quantum thermal bath has five input parameters, namely, target quantum temperature, temperature damping parameter, a random seed, upper cutoff frequency and the number of frequency bins. To make the results comparable with the results from classical molecular dynamics simulations, we set the temperature damping to be 0.3 ps (300 fs), the same one used in the purely classical simulations. The upper cutoff frequency was set to be three times the Debye frequency of silicon, due to the fact that SW potential overestimated the Debye frequency of silicon [66]. The number of frequency bins was 100 in all QTB simulations. When using the thermostat, a Berendsen barostat with the settings in section II.1 was also applied to the system.
To check whether these settings enable the compensation of low temperature quantum effect, the heat capacity of pure silicon crystal was calculated by differentiating the mean total energy with respect to temperature (an approach suggested by Dammak et al [38]), and the results was compared with an experiment [67]. The calculation results from two different thermostats, QTB thermostat and Berendsen thermostat, are compared with each other.
[…]
With the help of the quantum thermal bath, the annealing behavior of the system at very low temperatures can be investigated directly in molecular dynamics simulations. The five-stage simulation and analysis procedure in section II.1 was followed, and quantum thermal bath was used in stage(ii) for the annealing simulations.
Figure 2: Heat capacity of silicon crystal, calculated with the quantum thermal bath (QTB) and the classical Berendsen thermostat. Since the Berendsen thermostat and molecular dynamics describes only classical atom motion, its heat capacity prediction corresponds to the value given by the equipartition theorem, which does not have a temperature dependence [68]. The results are compared against experimental values (red plus signs).
III Results and discussions
@@ III.1 Stage (i) Damage production in cascade simulations
An example of a cascade event and the damage produced is shown in a cross-sectional plot in Figure 1. The cross-section shows a very typical damage production event, analyzed in detail in many previous works [69, 27, 57]. As noted above, 5 keV is above the sub-cascade threshold breakdown, so the damage is split into separate damage production regions. 1000 different 5 keV cascades were simulated, and the final atom coordinates from these were used as inputs for annealing runs, quenching and energy release analysis.
@@ III.2 Stage (ii) Annealing simulations
For the classical molecular dynamics simulations, annealing simulations were performed at several temperatures in the range 20 - 800 K for some of the cells, and at 200 K for all the cells. For the Tersoff potential, annealing was done at 300 K for all the cells. Since the Tersoff potential has a higher melting point than the Stillinger-Weber potential and the experimental one, 300 K corresponds to roughly the same homologous temperature as 200 K for the Stillinger-Weber potential. For simulations with QTB thermostat, the annealing simulations of 10 ns were conducted at 0.05 K, 10 K, 25 K, 50 K, 100 K, 150 K, 200 K and 300 K. At each temperature, at least 200 annealing simulations with different initial damage were carried out, that is to say, using different final states from different cascade simulations as input structures. Again, the number of annealing cases at each temperature is shown in Table 1.
Results for the number of defects in the cell as a function of time are shown in Figure 3 (a) for a test 600 K annealing run via classical molecular dynamics. The potential energy has huge fluctuations because a thermodynamic system with a small number of atoms $N$ has natural fluctuations of the kinetic energy around the average of the order of $\sqrt{N}/N$ [68]. Since the total energy remains the same on short time scales, the fluctuations in kinetic energy are also reflected in the potential energy. The fluctations make it very difficult to obtain a quantitative analysis of the potential energy release in a single annealing event. This was the reason why we must use the quenching scheme to get rid of thermal fluctuations.
@@
Figure 3: a) Annealing of damage produced in a single 5 keV cascade in Si at 600 K. The data shows that both the average potential energy and the number of Wigner-Seitz defect pairs decreases with time. b) Same as subfigure (a) but analyzed after each time step was quenched to 0 K. The inset shows the atomic positions in one disordered pocket before and after the annealing event occuring at $t=1.236$ ns. c) Potential energy release per 2 ps intervals from the same cascade. The inset shows that annealing at well-separated times keep occurring also at longer times and even at the cryogenic temperature of 100 K.
@@ III.3 Stage (iii) Quenching runs
Results of the potential energy and number of WS defect pairs after the quenching scheme (see section II) are shown in Figure 3 (b), for the exact same annealing run as shown in part 3 (a). This graph shows in a much clearer way that after the first about 10 ps, the annealing proceeds stepwise, with clear energy release events occuring every now and then, separated by (on the atomic scale) long times when essentially nothing else than thermal vibrations of atoms around their stable or metastable local equilibrium positions happens in the system. This phenomenon agrees very well with the observations by Caturla et al [70].
Figure 4 demonstrates the potential energy curves and the number of Wigner-Seitz defect pairs together. From this figure, one can easily see that the energy release correlates very well in time with decreases of the number of Wigner-Seitz defects, confirming that the potential energy release is indeed associated with partial annealing of damage in the amorphous pockets, especially the annealing on the amorphous-crystalline interface, where atoms from the amorphous side can rearrange themselves to become crystalline and consequently release energy.
We note that energy release events may occur either inside the amorphous pockets or at the interface between the crystalline and amorphous phase. There are several reports showing that fully amorphous materials can also undergo relaxation, for example, through the correlated atom motion [71, 72, 73, 74]. Such relaxations would not, however, on average reduce the number of Wigner-Seitz defect pairs, because the atomic structure there is still amorphous after the occurence of energy release events. In our simulations, the energy reduction steps in potential energy correlate almost always in time with a reduction in the number of Wigner-Seitz defect pairs (Figure 3 b and Figure 4). This indicates that the majority of the large energy release events occur at the interface, which is also observed when the a few selected annealing cases were analyzed in detail (cf. section V).
From the potential energy release, one can also obtain the magnitude of the energy released in each annealing event by comparing the total potential energies of the system before and after each 2 ps annealing interval. This is shown in Figure 3 (c) for the same annealing run as in parts (a) and (b) in Figure 3. The graph shows that the energy release varies from an annealing event to another.
@@
We ran similar annealing runs for a few cascades in the temperature range 20 – 800 K. Results for exemplary cases in Figure 4 show that, as expected, at the highest temperatures, very efficient annealing of the damage is observed. This observation in general agrees with the simulation [26] and the experimental [75, 76, 77, 78] reports showing that it is very difficult to have silicon amorphized by irradiation at temperatures above about 650 K.
However, we also observe that even at the much lower temperatures, down to 100 K, damage annealing is observed throughout the 3 ns simulations, even in the cases with the classical Berendsen thermostat. Also in some test annealing cases for 30 nanoseconds conducted by the Berendsen thermostat, annealing events kept occurring throughout the entire simulation time: an example of annealing occurring around 15 ns after the cascade event at 100 K is seen in the inset of Figure 3 c. Occasional annealing events are observed in the 50 K and 20 K simulations even in the classical simulations with no quantum thermal vibrations. An explanation to this surprising observation is given in section IV. These results are well in line with the previous work by Béland, Mousseau and co-workers who showed with a hybrid molecular dynamics/barrier search approach (“k-ART”) that annealing events can even occur on macroscopic, up to 1 s, time scales [35, 36, 37]. In the future, we are also going to work on the extension of time scales in our simulation work on defects in dark matter detector materials.
@@ III.4 Stage (iv) Statistics of energy release
To enable a comparison with the experimental observations of the energy dependence of the low energy excess, we made a statistics of the annealing data similar to that shown in Figure 4 but summing up all events over at least 50 independent cases (separate cascades as inputs) to obtain a statistically better picture of the annealing.
For classical molecular dynamics simulations, the results for the temperatures 100 - 500 K are shown in Figure 5 (a). Also included in this plot are results obtained with the Tersoff interatomic potential at 300 K. The results show that, remarkably, the energy release statistics is independent of temperature in the range 100 - 500 K within the statistical uncertainty. Moreover, the Tersoff interatomic potential gives results practically identical to those of the Stillinger-Weber potential. Since Tersoff and Stillinger-Weber potential are developed completely independent of each other, having different functional forms, and being fitted to different properties [52, 49], this gives high confidence that the obtained results are not an interatomic potential artifact.
For simulations with QTB themostat, the results at various temperatures are shown in Figure 5 (b). It is very clear that the energy release statistics are independent of temperature, since the magnitudes of the slopes are hardly related to the changes in temperature.
In both Figure 5 (a) and Figure 5 (b), the data are clearly close to linear on the log-lin plot, showing that the energy release statistics follows an exponential dependence with energy release magnitudes, in agreement with the experimental observations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. As expected from the linear dependence of the plots, the data can be fit very well with the functional form:
$f(E)=A\exp(-\alpha E)$ (3)
@@
a) Refer to caption
b) Refer to caption
Figure 5: Statistics of the magnitude of the energy release events during annealing runs at different temperatures. To enable easy comparison of the slopes, the curves at different temperatures have been scaled to having a value of 1 at 1 eV. a) Classical molecular dynamics, b) Molecular dynamics with the quantum thermal bath
Since the slope coefficient $\alpha$ is barely related to temperature, we can make an attempt to gather together all the energy release events at all the temperature levels (0.05 K, 10 K, 25 K, 50 K, 100 K, 150 K, 200 K, 300 K) in simulations with QTB thermostat, and plot a combined histogram of the energy release statistics of all the annealing events. This combined histogram is Figure 6. In this figure, 100 histogram bins are used, and the energy release spectrum can be fitted by exponential functions. However, it is worth noticing that we can obtain two slope coefficients by choosing the energy release range for fitting. The first slope $\alpha_{1}$ is calculated by fitting the entire histogram, while the second slope $\alpha_{2}$ is obtained by only fitting the energy range larger than 4.0 eV. We can clearly see that $\alpha_{2}$ is significantly smaller than $\alpha_{1}$. Compared with $\alpha_{1}$, we attach much more importance to $\alpha_{2}$, which in terms of energy is closer to experimental energy release spectra and can consequently be compared against experimental results.
Figure 6: The histogram of all the energy release events from all temperature levels. The energy release events from annealing simulations at 0.05 K, 10 K, 25 K, 50 K, 100 K, 150 K, 200 K, 300 K are gathered together to obtain an improved energy release statistics. In total, 28222 events are collected.
The simulated and experimental data from a silicon detector are compared to each other in Figure 7. The $y$ axis scale is in arbitrary units, since in experiments the absolute values of the signal depends on factors such as detector size, detection efficiency, concentration of radioactive impurities creating the initial damage, and damage recoil spectrum, all of which are not fully well known. However, the functional dependence of the signal on energy can be used for a direct comparison between simulation and experiments.
@@
The Figure 7 shows a very good agreement: both the simulation and low-energy tail of the experimental data clearly follow an exponential trend with a very similar slope. The slope obtained from a fit to the simulation data at 0.05 K (QTB thermostat) was $\alpha_{\rm sim}=0.58\pm 0.05$ 1/eV, which is in perfect agreement with a fit of the same exponential law to the low-energy tail of the experimental data [81] $\alpha_{\rm exp}=0.57\pm 0.01$ 1/eV (to enable direct comparison, this value was obtained by fitting the experimental data with the same software as used to fit the simulated ones). We emphasize that no fitting of the simulations to the experiments has been used for obtaining this outstanding agreement. Also, the slope coefficient in our combined histogram (Figure 6) is around 0.53 in the high energy regime, being very close to the experimental value. Based on very similar slope values, Figure 5, Figure 6 and Figure 7 together can indicate that the energy release events brought about by the defect evolution in detector materials can possibly be the source of the low energy excess signals in dark matter detectors.
Table 2: Slope coefficients $\alpha$ of the exponential dependence $\exp(-\alpha E)$ of the energy release $E$ from simulations and experiments on Si. The simulation results obtained with the classical Berendsen thermostat are marked “Berendsen”. and the ones with the quantum thermal bath “QTB”.
@@
Simulation 500 Berendsen 0.49 $\pm$ 0.03
(current work) 400 Berendsen 0.49 $\pm$ 0.03
300 Berendsen 0.53 $\pm$ 0.03
300 QTB 0.52 $\pm$ 0.01
200 Berendsen 0.52 $\pm$ 0.02
200 QTB 0.51 $\pm$ 0.02
100 Berendsen 0.60 $\pm$ 0.05
100 QTB 0.58 $\pm$ 0.07
50 QTB 0.52 $\pm$ 0.06
25 QTB 0.49 $\pm$ 0.04
10 QTB 0.50 $\pm$ 0.06
0.05 QTB 0.58 $\pm$ 0.05
@@ III.5 Stage (v) Analysis of annealing time scales
We finally analyze the time dependence of the annealing events such as those shown in Figure 3 and 4. To be able to roughly estimate the annealing time scales as a function of temperature and extrapolate to even lower temperatures, we analyzed the time dependence of the annealing curves.
@@
Since the number of disordered pockets in a single event is small, the annealing curves of single events have huge statistical fluctuations. This is illustrated in Figure 8. To obtain a statistically reliable analysis, we carried out an analysis of at least 200 different cascade events at each temperature level in the temperature range 0.05 – 300 K. To be able to compare cases with different numbers of initial defects and stored energy in them, the analysis was done for the average relative fraction of the excessive potential energy remaining at time $t$ compared to the initial excessive potential energy at $t=0$. The fraction $f$ was calculated for each cascade event separately before averaging. The results of the averaged fraction analysis are shown in Figure 9 (a). As expected, time evolution of the fraction of remaining Wigner-Seitz defects showed a similar behaviour.
The averaged annealing curves show that the annealing behaviour can be divided in two stages: a rapid annealing occurring roughly during the first 0.5 ns after the damage event, and a subsequent much slower annealing after this. The first stage clearly is due to annealing of the most metastable defects, whereas the later annealing is an average over occasional well-separated events, corresponding to the late energy release events shown in Figure 3.
By comparing the individual cases shown in Figure 8, an additional important point is illustrated: the damage in some events, especially those with large damage pockets, can anneal much slower than the average energy release. On long time scales, annealing from such cases would dominate the residual recovery signal.
a) Refer to caption
b) Refer to caption
Figure 9: a) Average potential energy curves with the quantum thermal bath after quenching. The values are rescaled so that the initial value is one. In this way, the curves show the relative amount of excessive potential energy. b) Results of fits of the annealing time parameter $\tau$ to data on the time dependence of potential energy in annealing, being a function of the annealing temperature. Since in the classical simulations practically no annealing processes are observed below 100 K, it was not possible to provide an annealing time constant for the lowest temperatures in classical cases.
Table 3: Time parameter $\tau$ in the fitting $A\exp(-t/\tau)$ of the potential energy curves. The error bar is the uncertainty estimate of the parameter obtained from the least squares fitting procedure. The results obtained by classical molecular dynamics with Berendsen thermostat are marked “Berendsen” and the ones with the quantum thermal bath “QTB”.
@@
150 Berendsen 348.8 $\pm$ 1.4
150 QTB 172.53 $\pm$ 0.46
125 Berendsen 451.5 $\pm$ 1.8
100 Berendsen 1037.8 $\pm$ 3.1
100 QTB 388.0 $\pm$ 1.3
50 QTB 769.8 $\pm$ 2.5
25 QTB 1070.9 $\pm$ 2.4
10 QTB 1234.0 $\pm$ 3.3
0.05 QTB 1326.1 $\pm$ 3.7
The curves in Figure 9 were fitted against the exponential function $A\exp(-t/\tau)$. In the fitting practice, the initial 1 ns in the potential energy curves was truncated, and the rest of the curves were fitted. The fitting results and the comparison with the results from purely classical annealing are shown in Figure 9 (b) and Table 3.
We note that in previous works [36], a roughly logarithmic time dependence, namely, a relaxation being linear with log(time) in annealing, was reported for damage in Fe. In the current case, we found that at room temperature, the time dependence was indeed logarithmic, but at the lower temperatures it was not, at least up to the time scale of 10 ns studied here. We also note that due to limited numbers of annealing events at the lower temperatures, the statistics is limited and we cannot be certain that the time dependence on even longer times can be described with the function $\exp(-t/\tau)$ with a single time constant. However, the current fits to a slowly decaying time dependence are sufficient to show that the randomly timed annealing events do occur up to at least millisecond time scales after the initial picosecond time scale radiation event. Hence the annealing energy release events may appear uncorrelated to the initial radiation production event.
From these results, it is very clear that the low temperature quantum effect can facilitate annealing significantly. This is because the quantum zero point energy brought by the quantum bath may trigger the atomic rearrangements as they only need to overcome barriers around $\lesssim 0.1$ eV (cf. Figs. 10 (b) and 11 (b)), which are comparable to the zero point energy $E_{\rm zp}\sim 0.06$ eV, using the Debye model estimate $E_{\rm zp}=9/8k_{B}\Theta_{D}$ and Debye temperature of $\Theta_{D}\approx 630$ K for silicon [64, 83]. At the lowest temperatures, these energy barriers are not accessible for the thermal vibrations in simulations with classical thermostats, since the classical thermal energy is $k_{B}T\ll 0.1$ eV when $T\lesssim 100$ K.
The simulation at 0.05 K with QTB thermostat shows that the annealing time parameter $\tau$ has a magnitude of microseconds at the operation temperature of the detectors. The fact that at 10 K and 0.05 K time constants are almost the same indicates that at temperatures $\lesssim 10$ K, the triggering of annealing effects is dominated by the quantum effect.
The exponential decay of annealing effects could be detectable if the time resolution of the experiments is shorter than the time constants $\tau$. In contrast, if time resolution is much longer than this, the energy release events would seem to appear at random times with average constant rate, which would be determined by the concentration of radioactive impurities that produce the initial damage.